Quantity of
Product A
Total Utility Marginal Utility Quantity of
Product B
Total Utility Marginal Utility

1

80

80 1 30 30
2 120 40 2 46 16
3 150 30 3 61 15
4 165

15

4 75 14
5 172 7 5 88 13
6 178 6 6 100 12
7 182 5 7 111 11
8 185 3 8 121 10

The table above refers to Elizabeth's utility schedule of Product A and Product B. The equilibrium price of Product A and Product B is $1 and $5 respectively. If Elizabeth has only $11 to spend on both goods, how many units of Product A will she consume if her goal is to maximize her utility?
Answer #1

Quantity of Product A MU of A/ Price of A Quantity of Total Product B Utility 30 Marginal Utility 30 MU of B / Price of B 80

Optimal condition is attained where MU of A / Price of A = MU of B / Price of B and budget constraint is satisfied.

MU of A / Price of A = MU of B / Price of B = 6 at 6 units of A and 1 unit of B

Budget line: PA.A + PB.B = Income

1A + 5B = 11

A + 5B = 11

Total cost of consuming 6 units of A and 1 unit of B = 1 x 6 + 5 x 1 = 11 = Income

So, Elizabeth will consume 6 units of A and 1 unit of B to maximize her utility.